# Design of low-energy transfers in cislunar space using sequences of lobe dynamics

### v1

@@ 5.1 Earth–Moon transfer design
To build a graph for this problem, the nodes corresponding to the start points, centroids of the effective lobes, and goal points are connected by edges, as shown in Fig. 21. The arrows indicate possible transfer paths, and a spacecraft can move from any selected effective lobe sequence to the LLO. Transfer arcs and their weights are determined by the targeting strategy shown in Fig. 17. The $|\Delta\bm{V}_{\mathrm{E}}|$ and $|\Delta\bm{V}_{\mathrm{M}}|$ are added to the cost function $J$ in Eq. (22) when the total transfer cost is evaluated. The weight threshold $w^{\ast}$ is set to $400$ m/s based on the graph structure in this case.
@@ 5.3 Role of lobe dynamics with respect to WSB
6 Extension of the obtained results into the BCR4BP
@@ 6.1 Optimization of the Earth–Moon transfer trajectory
The obtained results are summarized in Table 2. This table uses a “-” to denote that the optimization with the corresponding $\theta_{s}^{\ast}$ does not converge after $50000$ iterations. All the converged solutions have a transfer time similar to that of the initial guess ($191.8994$ days), but the total $\Delta V$ varies significantly. The best solution in this table is the one with $\theta_{s}^{\ast}=0$. Figure 26 illustrates the optimal trajectory in the BCR4BP with $\theta_{s}^{\ast}=0$. The dashed line in this figure represents the optimal trajectory in the CR3BP (Fig. 23), while the solid line indicates the optimal trajectory in the BCR4BP. The triangle and star in this figure correspond to the start and goal points, respectively. Similar to Fig. 23, blue arrows indicate $\Delta\bm{V}$ vectors, and the length of these arrows corresponds to the relative magnitude of the $\Delta\bm{V}$s. Unlike the optimal transfer in the CR3BP, the majority of the total $\Delta V$ is contributed solely by the two largest $\Delta\bm{V}$s. Specifically, $|\Delta\bm{V}_{\mathrm{E}}|=3120.8662$ m/s is applied to depart the LEO, and $|\Delta\bm{V}_{\mathrm{M}}|=638.5176$ m/s is applied to arrive at the LLO. The other small $\Delta\bm{V}$s satisfying $|\Delta\bm{V}|<21$ m/s are used to adjust the transfer trajectory.
@@
6.2 Comparison of the obtained results with literature

### v2

@@ 5.1 Earth–Moon transfer design
To build a graph for this problem, the nodes corresponding to the start points, centroids of the effective lobes, and goal points are connected by edges, as shown in Fig. 24. The arrows indicate possible transfer paths, and a spacecraft can move from any selected effective lobe sequence to the LLO. Transfer arcs and their weights are determined by the targeting strategy shown in Fig. 18. The $|\Delta\bm{V}_{\mathrm{E}}|$ and $|\Delta\bm{V}_{\mathrm{M}}|$ are added to the cost function $J$ in Eq. (22) when the total transfer cost is evaluated. The threshold weight $w^{\ast}$ is set to $400$ m/s based on the graph structure in this case.
@@ 5.3 Role of lobe dynamics with respect to WSB
6 Extension of the CR3BP results into the BCR4BP
@@ 6.1 Optimization of the Earth–Moon transfer trajectory
The obtained results are summarized in Table 2. This table uses a “-” to denote that the optimization with the corresponding $\theta_{s}^{\ast}$ does not converge after $50000$ iterations. The failure of convergence for some values of $\theta_{s}^{\ast}$ is due to the sensitivity of the initial guess to $\theta_{s}^{\ast}$. These results imply that the convergence may be improved by gradually adding the perturbation of the Sun. All the converged solutions have a transfer time similar to that of the initial guess ($191.8994$ days), but the total $\Delta V$ varies significantly. The best solution in this table is the one with $\theta_{s}^{\ast}=0$. Figure 29 illustrates the optimal trajectory in the BCR4BP with $\theta_{s}^{\ast}=0$. The dashed line in this figure represents the optimal trajectory in the CR3BP (Fig. 26), while the solid line indicates the optimal trajectory in the BCR4BP. The triangle and star in this figure correspond to the start and goal points, respectively. Similar to Fig. 26, blue arrows indicate $\Delta\bm{V}$ vectors, and the length of these arrows corresponds to the relative magnitude of the $\Delta\bm{V}$s. Unlike the optimal transfer in the CR3BP, the majority of the total $\Delta V$ is contributed solely by the two largest $\Delta\bm{V}$s. Specifically, $|\Delta\bm{V}_{\mathrm{E}}|=3120.8662$ m/s is applied to depart the LEO, and $|\Delta\bm{V}_{\mathrm{M}}|=638.5176$ m/s is applied to arrive at the LLO. The other small $\Delta\bm{V}$s satisfying $|\Delta\bm{V}|<21$ m/s are used to adjust the transfer trajectory.
6.2 Comparison of the results with the literature